1. Imagine that you work in a World Bank office that is in charge of making recommendations to increase the growth of the economy in Nigeria. According to the information you have available, the population of Nigeria is growing at an annual rate of 5%, capital depreciates at 3% annually, the savings rate is 10% and Nigeria's economy requires an additional unit of capital to produce an additional unit of product. Your direct boss tells you that in the office they use the Harrod-Domar model.

1. (a) According to the Harrod-Domar model, what is the growth rate of the economy in per-capita terms?
Answer #1

in Harrod Domar Model

At eqm,

(S/ θ) = g + n + d

d : depreciation rate = 3%

g : growth rate of Economy

n: population growth rate =5%

S : Saving rate = 10%

θ : capital - output ratio = K/Y = 1

So as per the formula

10 = g + 5 + 3

so g* = 2%